# Superpose Singular Features for Model Merging

### v1

@@ Abstract
Model merging is a critical technique for combining the capabilities of multiple fine-tuned models without requiring additional training. While existing methods treat parameters as vectors, they overlook the intrinsic structure of linear transformation matrices - the core components that comprise the majority of model parameters. These matrices are fundamental to neural networks, mapping input representations to output features through linear combinations. Motivated by the linear representation hypothesis, we introduce task matrix and propose to Superpose Features from Task Matrix (SFTM), a novel approach that superposes features from individual task models into a merged model. SFTM employs singular value decomposition to identify feature bases of linear transformation matrices and solves a linear system to optimally combine them while preserving input-output mappings from individual task models. Extensive experiments on vision transformers and language models demonstrate that SFTM consistently outperforms existing methods, achieving superior performance and enhanced out-of-distribution generalization.
@@ 1 Introduction
Pre-trained models have become a cornerstone of modern deep learning, providing a powerful foundation for a wide range of tasks (Zhuang et al., 2020; Bommasani et al., 2021). To serve different tasks, fine-tuning the pre-trained models separately on different tasks is a common practice (Kenton and Toutanova, 2019; Hu et al., 2021). However, maintaining multiple fine-tuned models for different tasks can be cumbersome and resource-intensive, especially as the number of tasks and models continues to grow. Multitask learning (Sanh et al., 2021; Raffel et al., 2020b) is a promising approach to address this issue by training a single model on multiple tasks simultaneously, but it often requires a large amount of data and computational resources to achieve optimal performance (Pilehvar and Camacho-Collados, 2018; Phang et al., 2018; Fifty et al., 2021).
Recently, model merging (Matena and Raffel, 2022; Jin et al., 2022; Ilharco et al., 2022a; Yadav et al., 2024) offers a solution to this problem by consolidating multiple models into a single model that can perform well on multiple tasks effectively. This approach eliminates the need for maintaining multiple separate models, reduces storage and computational requirements, and potentially enables knowledge transfer between tasks. A foundational approach to model merging is parameter averaging (Choshen et al., 2022; Ilharco et al., 2022b; Wortsman et al., 2022), which simply takes the mean of corresponding parameters across models. This evolved into weighted averaging methods like Fisher Merging (Matena and Raffel, 2022) and RegMean (Jin et al., 2022) that assign different weights to parameters based on their importance. A key advancement came with the introduction of task vectors (Ilharco et al., 2022a) which represent the difference between fine-tuned and pre-trained weights and enable lightweight cross-task generalization. Various techniques (Huang et al., 2023; Yang et al., 2023; Yadav et al., 2024; Yu et al., 2024; Du et al., 2024) were subsequently developed to handle task conflicts by resetting redundant parameters, resolving sign conflicts, dropping and rescaling parameters, and managing parameter competition between tasks. However, these methods often rely on heuristic rules for merging parameters while overlooking the intrinsic structure of parameters.
Our method is motivated by the linear representation hypothesis (Mikolov et al., 2013; Arora et al., 2016; Olah et al., 2020). The linear representation hypothesis states that the representations of deep neural networks can be decomposed into combinations of feature vectors. This linearity enables neural networks to represent complex concepts through these combinations. Linear transformations (e.g., in linear and convolution layers) play a crucial role in this process, as they can activate specific features by measuring the similarity between a representation and a feature vector (via the inner product with the row vectors of the transformation matrix) and extract new features through linear combinations (via the weighted sum of the column vectors of the transformation matrix). These properties have been leveraged in recent years to interpret LLMs (Elhage et al., 2022; Bricken et al., 2023; Templeton et al., 2024) and design more efficient fine-tuning techniques (Hu et al., 2021; Zhang et al., 2023b). Consequently, in model merging, it is essential that the linear transformations in the merged model preserve the input-output mappings of those in the individual models for effective feature activation and extraction.
In this paper, we propose a method for merging linear transformation matrices while preserving the input-output mappings of individual models’ linear transformations. Our approach leverages Singular Value Decomposition (SVD) to merge matrices, treating left singular vectors as basis for output representations and right singular vectors as basis for input representations. The goal is to ensure the merged matrices preserve the directions of output features from original transformations when applied to corresponding input features across tasks. We show this problem reduces to solving a system of linear equations, enabling model merging in singular decomposition space while maintaining feature directions from various tasks. We validate our method on multiple benchmark datasets spanning vision and natural language processing tasks. The results demonstrate that our method consistently outperforms existing model merging techniques.
@@
Throughout this paper, we use bold uppercase letters (e.g., $\mathbf{X}$) to denote matrices, bold lowercase letters (e.g., $\mathbf{x}$) to denote vectors, and regular lowercase letters (e.g., $x$) to denote scalars. For a matrix $\mathbf{X}$, its transpose is denoted as $\mathbf{X}^{\top}$. We denote task $i$ as $\mathsf{T}_{i}$ and its corresponding model parameters as $\theta_{i}$. For linear transformations, a linear transformation matrix in the model trained on task $i$ is represented as $\mathbf{P}_{i}$. The subscript ${}_{\text{pre}}$ indicates pre-trained parameters or matrices - for example, $\theta_{\text{pre}}$ and $\mathbf{P}_{\text{pre}}$ denote the parameters and linear transformation matrices of the pre-trained model, respectively.
@@ 2.1 Model Merging of Fine-tuned Models
Model merging is a technique that combines multiple models into a single model to enhance performance or enable the model to perform multiple tasks. Previous studies have shown that averaging the weights of multiple models fine-tuned from the same pre-trained initialization is a promising approach for model merging. Fisher Merging (Matena and Raffel, 2022) advances beyond simple averaging by utilizing the Fisher information matrix to assess the importance of individual parameters, which are then weighted accordingly during the merging process. Similarly, RegMean (Jin et al., 2022) forms a linear regression problem with extra data for each layer and offers a closed-form solution for the merged model’s parameters by solving the regression problem.
Beyond parameter averaging, Task Arithmetic (Ilharco et al., 2022a) introduces task vectors and adding the task vectors of individual tasks to merge model, demonstrating their effectiveness and lightweight nature in facilitating cross-task generalization. Building on this concept, PEM Composition (Zhang et al., 2023a) extends the task arithmetic framework to merge LoRA (Hu et al., 2021), while Ties-Merging (Yadav et al., 2024) addresses task conflicts by resetting redundant parameters and resolving sign conflicts. These methods, however, use a single merging coefficient across all task vectors, which limits their flexibility. In contrast, Lorahub (Huang et al., 2023) and AdaMerging (Yang et al., 2023) use different coefficients for enhanced adaptability. Lorahub’s performance is limited as it only searches for coefficients at the task level, while AdaMerging requires complex training and unlabeled test datasets, making it applicable solely to classification problems. DARE (Yu et al., 2024) proposes drop and rescale as preprocessing steps when merging fine-tuned LLMs. PCB-Merging (Du et al., 2024) is a lightweight, training-free technique for model merging that balances parameter competition by intra-balancing parameter significance within tasks and inter-balancing parameter similarities across tasks, effectively enhancing performance across various scenarios and domains.
@@ 2.2 Linear Representation Hypothesis
The linear representation hypothesis states that neural networks encode information by summing up "feature vectors" (Mikolov et al., 2013; Arora et al., 2016; Olah et al., 2020), i.e., a layer of a network represents a set of features as a weighted sum of task-associated vectors. This hypothesis has been observed in various models, including word embeddings (Mikolov et al., 2013; Conneau et al., 2017), sentence embeddings (Bowman et al., 2015), Transformer language models (Meng et al., 2022; Hendel et al., 2023), and vision-language models (Trager et al., 2023; Perera et al., 2023). The hypothesis has been explioted in various fields, especially in probing (Alain and Bengio, 2018; Belinkov, 2022) and interpretability (nostalgebraist, 2020; Elhage et al., 2022; Bricken et al., 2023; Gao et al., 2024).
Especially, the linear representation hypothesis is prominently featured in recent works on mechanistic interpretablity of language models (Olsson et al., 2022; Elhage et al., 2022; Bricken et al., 2023; Templeton et al., 2024). In mechanistic interpretability, models are understood by decomposing them into interpretable components and understanding how these components interact. The linear representation hypothesis suggests that important features in neural networks are often represented through linear combinations of neuron activations, rather than complex nonlinear transformations. This hypothesis has gained significant empirical support through studies of language models. For instance, (Elhage et al., 2022) demonstrated a toy model that learned to resconstruct its input through linear combinations of its neurons. Furthermore, Bricken et al. (2023) showed that individual neurons in a language model can encode semantically meaningful features in a largely linear fashion, which is later sclaed to large language models (Templeton et al., 2024).
@@ 3.1 Problem Formulation
We start with a set of tasks $\{\mathsf{T}_{1},\ldots,\mathsf{T}_{T}\}$ and various pre-trained models. The objective is to fine-tune these models either by updating all parameters or using parameter-efficient fine-tuning (PEFT) methods. The goal of model merging is to combine multiple fine-tuned models $\{\theta_{1},\ldots,\theta_{n}\}$ into a single model $\theta_{M}$ that can perform all tasks $\{\mathsf{T}_{1},\ldots,\mathsf{T}_{T}\}$ effectively without requiring access to training data. While existing methods treat model parameters as vectors to be merged, we focus on the crucial role of linear transformations, which comprise the majority of parameters and computations in neural networks. These transformations are essential as they project data into spaces where important features become more apparent, making information processing more efficient. To formally analyze these transformations during merging, we introduce the Task Matrix in Definition 3.1.
Definition 3.1 (Task Matrix).
The task matrix is defined by $\mathbf{M}_{i}=\mathbf{P}_{i}-\mathbf{P}_{\text{pre}}\in\mathbb{R}^{m\times n},$ where $\mathbf{P}_{i}$ is a linear transformation matrix for model of task $\mathsf{T}_{i}$, $\mathbf{P}_{\text{pre}}$ is the linear transformation matrix for the pre-trained model.
Remark 3.2.
Existing merging methods based on task vectors neglect the internal structure of linear transformation in neural networks, and merge parameters as vector, i.e., $\text{vec}(\mathbf{P}_{j})=\text{vec}(\mathbf{P}_{\text{pre}})+\gamma\text{vec}(\mathbf{M})$. These methods employ various heuristic rules to merge parameters, which requires many hyperparameter tuning.
@@
To merge models, preserving the input-output mappings of linear transformations is essential for maintaining model capabilities. Our method focuses specifically on merging linear transformations in a way that retains their key input-output mappings through singular value decomposition, rather than treating all parameters equally as in previous approaches. Specifically, we focus on merging task matrices $\mathbf{M}_{i},i=1\ldots T$ into a single merged matrix $\mathbf{M}\in\mathbb{R}^{m\times n}$. For a input feature $\mathbf{r}\in\mathbb{R}^{n}$, we aim to ensure that the output of the merged model $\mathbf{M}\mathbf{r}$ maintains the same directional features as the outputs $\mathbf{M}_{i}\mathbf{r}$ from the original task-specific models, i.e., the merged model preserves the input-output mappings of the individual models. The final merged model is then constructed by adding the merged task matrix $\mathbf{M}$ to the pre-trained model with a scaling factor $\gamma$: $\mathbf{P}_{j}=\mathbf{P}_{\text{pre}}+\gamma\mathbf{M}$.
3.2 Merge Task Matrices to Superpose Singular Features
Identify Basis for Preserving Mappings
@@ 3.2 Merge Task Matrices to Superpose Singular Features
To keep the directional features of the task matrices $\mathbf{M}_{i}\mathbf{r}$ during merging, we first need to identify some basis features that is representative for the input and output representations. Singular value decomposition (SVD) of task matrices provides a natural way to obtain the basis features. Given the decomposition $\mathbf{M}_{i}=\sum_{k=1}^{r_{i}}\sigma_{i}^{(k)}\mathbf{u}_{i}^{(k)}\mathbf{v}_{i}^{(k)\top}$ where $\sigma_{i}^{(k)}$ is the $k$-th singular value, $\mathbf{u}_{i}^{(k)}\in\mathbb{R}^{m\times 1}$ and $\mathbf{v}_{i}^{(k)}\in\mathbb{R}^{n\times 1}$ are the $k$-th left and right singular vectors of $\mathbf{M}_{i}$, respectively. Therefore, the right singular vectors $\mathbf{v}_{i}^{(k)}$ form a basis for input representations, while the left singular vectors $\mathbf{u}_{i}^{(k)}$ form a basis for output representations. We named these basis features as singular features in our paper.
Except form the basis, singular features have several other properties that make them suitable for preserving input-output mappings during merging. First, the orthogonality of singular vectors within each task minimizes interference between features, enabling linear superposition of features in accordance with the linear representation hypothesis. Second, the decomposition reveals the model’s information processing mechanism - using inner products to detect input features and weighted sums to construct outputs. We leave the details of the good properties of singular features to the appendix.
Merge Task Matrices in Singular Space
Since we have the singular vectors as the basis of the input and output representations for linear transformation $\mathbf{M}_{i}$, the merged matrix $\mathbf{M}$ should be able to keep the directions of output features from original transformations when applied to corresponding input features across tasks. To accomplish this, we can merge the task matrices $\mathbf{M}_{i}$ in their singular spaces:
$\mathbf{M}=\sum\nolimits_{i=1}^{T}\sum\nolimits_{k=1}^{r_{i}}\alpha_{i}^{(k)}\sigma_{i}^{(k)}\mathbf{u}_{i}^{(k)}\mathbf{v}_{i}^{(k)\top},$ (1)
where $\alpha_{i}^{(k)}$ are the merging weights. To preserve feature directions during merging, we require that when the merged matrix $\mathbf{M}$ transforms an input representation, the output should maintain the same directional features as those produced by the individual task matrices $\mathbf{M}_{i}$. This leads to the following objective:
$\mathbf{u}_{i}^{(k)},\mathbf{M}\mathbf{v}_{i}^{(k)}\right>=\left<\mathbf{u}_{i}^{(k)},\mathbf{M}_{i}\mathbf{v}_{i}^{(k)}\right>,\quad\forall i,j,k.$ (2)
[…]
Solving Linear Systems to Merge Task Matrices
To solve (2), we rewrite it as a linear system for all tasks in $\{\mathsf{T}_{1},\ldots,\mathsf{T}_{T}\}$:
$\begin{split}&\mathbf{u}_{i}^{(k)\top}\!\!\left(\sum_{i^{\prime}=1}^{T}\sum_{k^{\prime}=1}^{r_{i}}\!\alpha_{i^{\prime}}^{(k^{\prime})}\sigma_{i^{\prime}}^{(k^{\prime})}\mathbf{u}_{i^{\prime}}^{(k^{\prime})}\mathbf{v}_{i^{\prime}}^{(k^{\prime})\top}\!\!\right)\!\mathbf{v}_{i}^{(k)}\!=\!\sigma_{i}^{(k)}\\ \Leftrightarrow&\sum_{i^{\prime}=1}^{T}\sum_{k^{\prime}=1}^{r_{i}}\!\!\sigma_{i^{\prime}}^{(k^{\prime})}\mathbf{u}_{i}^{(k)\top}\mathbf{u}_{i^{\prime}}^{(k^{\prime})}\mathbf{v}_{i^{\prime}}^{(k^{\prime})\top}\mathbf{v}_{i}^{(k)}\alpha_{i^{\prime}}^{(k^{\prime})}\!=\!\sigma_{i}^{(k)}.\end{split}$ (3)
For various tasks $\mathsf{T}_{i}$ and singular values, (3) forms a linear system with $r=\sum_{i=1}^{T}r_{i}$ variables $\alpha_{i^{\prime}}^{(k^{\prime})}$ and $r$ equations. The coefficient matrix of this linear system is the element-wise product of three matrices:
  • $\mathbf{\Sigma}=\mathbf{1}\otimes[\sigma_{i^{\prime}}^{(k^{\prime})}]_{i^{\prime},k^{\prime}}\in\mathbb{R}^{r\times r}$, where $\mathbf{1}\in\mathbb{R}^{r\times 1}$, $\otimes$ is the Kronecker product, and $[\sigma_{i^{\prime}}^{(k^{\prime})}]_{i^{\prime},k^{\prime}}$ is a row vector of singular values;
  • $\mathbf{U}=[\mathbf{u}_{i}^{(k)\top}\mathbf{u}_{i^{\prime}}^{(k^{\prime})}]_{i,k,i^{\prime},k^{\prime}}\in\mathbb{R}^{r\times r}$, where enumerating over $i,k$ forms the row index and enumerating over $i^{\prime},k^{\prime}$ forms the column index;
  • $\mathbf{V}=[\mathbf{v}_{i^{\prime}}^{(k^{\prime})\top}\mathbf{v}_{i}^{(k)}]_{i^{\prime},k^{\prime},i,k}\in\mathbb{R}^{r\times r}$, where enumerating over $i,k$ forms the row index and enumerating over $i^{\prime},k^{\prime}$ forms the column index.
Then, the coefficient matrix of the linear system is $\mathbf{\Sigma}\circ\mathbf{U}\circ\mathbf{V}$, where $\circ$ is the element-wise product. The linear equation (3) can be written as
$\mathbf{\Sigma}\circ\mathbf{U}\circ\mathbf{V}\boldsymbol{\alpha}=\boldsymbol{\sigma},$ (4)

### v2

# Superpose Task-specific Features for Model Merging

@@ Abstract
Model merging enables powerful capabilities in neural networks without requiring additional training. In this paper, we introduce a novel perspective on model merging by leveraging the fundamental mechanisms of neural network representation. Our approach is motivated by the linear representation hypothesis, which states that neural networks encode information through linear combinations of feature vectors. We propose a method that superposes task-specific features from individual models into a merged model. Our approach specifically targets linear transformation matrices, which are crucial for feature activation and extraction in deep networks. By formulating the merging process as a linear system, we can preserve task-specific features from individual models and create merged models that effectively maintain multi-task capabilities compared to existing methods. Extensive experiments across diverse benchmarks and models demonstrate that our method outperforms existing techniques. Code is available at https://github.com/LARS-research/STF.
Superpose Task-specific Features for Model Merging
[…]
@@ 1 Introduction
(a) Task-specific feature preservation of merged methods
(b) Performance on T5 model merging
Figure 1: We measure the task-specific feature preservation of merged methods in (a) (refer to Appendix A), where smaller values indicate better preservation. The merging performance in (b) strongly correlates with task-specific feature preservation.
Because of the increasing resource demands of training models, model merging has emerged as an efficient approach to consolidate capabilities from specialist models while reducing storage and deployment costs. Various model merging methods have been proposed, including parameter averaging Choshen et al. (2022); Wortsman et al. (2022) and task vectors Ilharco et al. (2022); Du et al. (2024). However, these methods primarily focus on parameter-level operations and do not explicitly incorporate the fundamental working mechanisms of neural networks in their design.
We argue that a principled approach to model merging should be conditioned on how deep neural networks represent and process information. Therefore, to design our merging method, we draw upon the linear representation hypothesis Mikolov et al. (2013); Arora et al. (2016); Olah et al. (2020), which states that neural network representations can be decomposed into combinations of feature vectors. Recent works in mechanistic interpretability Elhage et al. (2022); Bricken et al. (2023); Templeton et al. (2024); Lindsey et al. (2025) validate the hypothesis and also reveal that these representations often contain features both related and unrelated to the model input. This phenomenon motivates our approach: linearly superposing features from individual models into the representation of the merged model can preserve task-specific capabilities.
Therefore, we propose a method that Superposes Task-specific Features(STF) from individual finetuned models. Our approach specifically targets linear transformation matrices, which comprise the majority of model parameters and form the foundation of the linear representation hypothesis. These matrices activate features through inner products with row vectors and extract new features through linear combinations via column vectors, making them crucial for detecting and processing information in deep neural networks. We design the merged linear transformation matrices to preserve the output features from individual models when processing the same input. By identifying task-specific features through singular value decomposition of task matrices, we formulate feature superposition as a linear system that derives the optimal merged task matrix. The final merged model is created by adding this merged matrix to the pre-trained model. Experiments verified the strong positive correlation between feature preservation and performance improvement(see Fig.˜1). We validate our method on multiple benchmark datasets and models. The results demonstrate that our method consistently outperforms existing model merging techniques.
@@
Throughout this paper, we use bold uppercase letters (e.g., $\mathbf{X}$) to denote matrices, bold lowercase letters (e.g., $\mathbf{x}$) to denote vectors, and regular lowercase letters (e.g., $x$) to denote scalars. For a matrix $\mathbf{X}$, its transpose is denoted as $\mathbf{X}^{\top}$. We denote task $i$ as $\mathsf{T}_{i}$ and its corresponding model parameters as $\theta_{i}$. For linear transformations, a linear transformation matrix in the model trained on task $i$ is represented as $\mathbf{P}_{i}$. The subscript $_{\text{pre}}$ indicates pre-trained parameters or matrices - for example, $\theta_{\text{pre}}$ and $\mathbf{P}_{\text{pre}}$ denote the parameters and linear transformation matrices of the pre-trained model, respectively. The notation $\mathbf{X}\circ\mathbf{Y}$ denotes the Hadamard product of matrices $\mathbf{X}$ and $\mathbf{Y}$.
@@ 2.1 Model Merging of Fine-tuned Models
Model merging is a technique that combines multiple models into a single model to enhance performance or enable the model to perform multiple tasks. Previous studies have shown that averaging the weights of multiple models fine-tuned from the same pre-trained initialization is a promising approach for model merging. Fisher Merging Matena and Raffel (2022) advances beyond simple averaging by utilizing the Fisher information matrix to assess the importance of individual parameters, which are then weighted accordingly during the merging process. Similarly, RegMean Jin et al. (2022) forms a linear regression problem with extra data for each layer and offers a closed-form solution for the merged model’s parameters by solving the regression problem.
Beyond parameter averaging, Task Arithmetic Ilharco et al. (2022) introduces task vectors and adding the task vectors of individual tasks to merge model, demonstrating their effectiveness and lightweight nature in facilitating cross-task generalization. Building on this concept, PEM Composition Zhang et al. (2023) extends the task arithmetic framework to merge LoRA Hu et al. (2021), while Ties-Merging Yadav et al. (2023) addresses task conflicts by resetting redundant parameters and resolving sign conflicts. These methods, however, use a single merging coefficient across all task vectors, which limits their flexibility. In contrast, Lorahub Huang et al. (2023) and AdaMerging Yang et al. (2023) use different coefficients for enhanced adaptability. Lorahub’s performance is limited as it only searches for coefficients at the task level, while AdaMerging requires complex training and unlabeled test datasets, making it applicable solely to classification problems. DARE Yu et al. (2024) proposes drop and rescale as preprocessing steps when merging fine-tuned LLMs. PCB-Merging Du et al. (2024) is a lightweight, training-free technique for model merging that balances parameter competition by intra-balancing parameter significance within tasks and inter-balancing parameter similarities across tasks, effectively enhancing performance across various scenarios.
Figure 2: Deep neural networks (a) linearly add both relevant and irrelevant features in hidden representations. This insight motivates our approach to model merging through linear superposition of task-specific features. The second row illustrates how STF preserves and combines essential features from individual models to create an effective merged method.
@@ 2.2 Linear Representation Hypothesis
The linear representation hypothesis states that neural networks encode information by summing up feature vectors Mikolov et al. (2013); Arora et al. (2016); Olah et al. (2020), i.e., a layer of a network represents a set of features as a weighted sum of task-associated vectors. This hypothesis has been observed in various models, including word embeddings Mikolov et al. (2013); Conneau et al. (2017), sentence embeddings Bowman et al. (2015), Transformer language models Meng et al. (2022); Hendel et al. (2023), and vision-language models Trager et al. (2023); Perera et al. (2023). The hypothesis has been exploited in various fields, especially in probing Alain and Bengio (2018); Belinkov (2022) and interpretability nostalgebraist (2020); Elhage et al. (2022); Bricken et al. (2023); Gao et al. (2024); Lindsey et al. (2025).
The linear representation hypothesis has been validated in recent research on mechanistic interpretability of language models Elhage et al. (2022); Bricken et al. (2023); Templeton et al. (2024), which showed that neural networks learn to represent meaningful features through linear combinations of neurons.
@@ 3.1 Problem Formulation
We start with a set of tasks $\{\mathsf{T}_{1},\ldots,\mathsf{T}_{T}\}$ and various pre-trained models. The objective is to fine-tune these models either by updating all parameters or using parameter-efficient fine-tuning (PEFT) methods. The goal of model merging is to combine multiple fine-tuned models into a single model that can perform all tasks $\{\mathsf{T}_{1},\ldots,\mathsf{T}_{T}\}$ effectively without requiring access to training data.
@@
To superpose task-specific features, we focus on merging the linear transformation matrices of the models, as these matrices are core components contributed to the linear representation hypothesis Elhage et al. (2022). In this paper, instead of merging linear transformation matrices directly, we merge the task matrices $\mathbf{M}_{i}=\mathbf{P}_{i}-\mathbf{P}_{\text{pre}}\in\mathbb{R}^{m\times n}$ where $\mathbf{P}_{i}$ is a linear transformation matrix for model of task $\mathsf{T}_{i}$, $\mathbf{P}_{\text{pre}}$ is the linear transformation matrix for the pre-trained model, as the task matrices contain the task-specific information compared to model parameters that contains both task-specific and pre-trained information. Specifically, we merge task matrices $\{\mathbf{M}_{i}\}$, i.e., into a single merged matrix $\mathbf{M}\in\mathbb{R}^{m\times n}$. For an input feature $\mathbf{x}\in\mathbb{R}^{n}$, we aim to ensure that the output of the merged model $\mathbf{M}\mathbf{x}$ maintains the features of the outputs $\mathbf{M}_{i}\mathbf{x}$ of the models for task $\mathsf{T}_{i}$. The final merged model is then constructed by adding the merged matrix $\mathbf{M}$ to the pre-trained model with a scaling factor $\gamma$: $\mathbf{P}=\mathbf{P}_{\text{pre}}+\gamma\mathbf{M}$.
3.2 Task-specific Feature Superposition
Identify Task-specific Features
@@ 3.2 Task-specific Feature Superposition
To merge the linear transformation matrices to linearly superpose task-specific features, we need to identify these features first. In this paper, we choose the singular vectors of the task matrices as the task-specific features, i.e., decomposing $\mathbf{M}_{i}$ as $\mathbf{M}_{i}=\sum_{k=1}^{r_{i}}\sigma_{i}^{(k)}\mathbf{u}_{i}^{(k)}\mathbf{v}_{i}^{(k)\top}$, where $\sigma_{i}^{(k)}$ is the $k$-th singular value, and $\mathbf{u}_{i}^{(k)}\in\mathbb{R}^{m}$ and $\mathbf{v}_{i}^{(k)}\in\mathbb{R}^{n}$ are the corresponding left and right singular vectors. This decomposition naturally provides task-specific features: the right singular vectors $\mathbf{v}_{i}^{(k)}$ form input features for the input $\mathbf{x}$, and $\sigma_{i}^{(k)}\mathbf{u}_{i}^{(k)}$ forms the task-specific feature for the output of task matrices.
Remark 1.
Singular vectors also reveal how the model processes information, i.e., the right singular vectors act as feature detectors through inner products with inputs, while the left singular vectors and singular values determine how detected features are combined into outputs. Refer to Appendix˜B for detailed analysis of singular vector.
Objective of Superposing Features
Given the input features $\mathbf{v}_{i}^{(k)}$ of task $i$, we want that after the features transformed by merged matrix $\mathbf{M}$, the output $\mathbf{M}\mathbf{v}_{i}^{(k)}$ contains the task-specific output feature $\sigma_{i}^{(k)}\mathbf{u}_{i}^{(k)}$. This leads to the following objective:
$\big<\sigma_{i}^{(k)}\mathbf{u}_{i}^{(k)},\underbrace{\mathbf{M}\mathbf{v}_{i}^{(k)}-\sigma_{i}^{(k)}\mathbf{u}_{i}^{(k)}}_{\text{Unrelated features}}\big>=0,$ (1)
where $\left$$\cdot$$\cdot\right$> is the inner product. In (1), we first obtain the unrelated features by subtracting the task-specific output feature $\sigma_{i}^{(k)}\mathbf{u}_{i}^{(k)}$ from the transformed feature $\mathbf{M}\mathbf{v}_{i}^{(k)}$. The inner product (1) should be zero, which means that the unrelated features are orthogonal to the task-specific output feature. This ensures that the merged matrix $\mathbf{M}$ preserves the task-specific feature $\sigma_{i}^{(k)}\mathbf{u}_{i}^{(k)}$ while allowing other features to be superposed linearly.
Remark 2 (Comparison against RegMean).
Ideally, we would fully preserve features by requiring $\mathbf{M}\mathbf{v}_{i}^{(k)}=\mathbf{M}_{i}\mathbf{v}_{i}^{(k)}$ for all tasks (similar to RegMean). However, due to the vast number of features, merged parameters cannot simultaneously maintain all features, as this would lead to an overdetermined system of equations. Instead, our approach, STF, superposes task-specific features rather than requiring exact feature preservation, resulting in increased representation efficiency and superior performance.
3.3 Merging by Solving Linear System
To merge the task matrices, we can see that the superposition objective (1) is a linear equation with the unknowns being the merging matrices $\mathbf{M}$. However, direct solving $\mathbf{M}$ is not feasible because of numerous unknowns and equations in the linear system consisting of (1). For efficient merging, we instead merge the task matrices $\mathbf{M}_{i}$ by decomposing them via SVD and merging their singular decomposition components:
$\mathbf{M}=\sum\nolimits_{i=1}^{T}\sum\nolimits_{k=1}^{r_{i}}\alpha_{i}^{(k)}\mathbf{u}_{i}^{(k)}\mathbf{v}_{i}^{(k)\top},$ (2)
where $\alpha_{i}^{(k)}$ are the merging weights. This reduces the number of unknowns and equations to $r=\sum_{i=1}^{T}r_{i}$, where $r_{i}$ is the rank of task matrix $\mathbf{M}_{i}$. By merging in the singular space, we can ensure that the resulting problem becomes a linear system of equations with the following theorem (proved in Appendix C).
Theorem 1.
Given the task matrices $\mathbf{M}_{i}$ with their SVD decompositions, the merging weights $\alpha_{i}^{(k)}$ can be obtained by solving the linear system:
$(\mathbf{U}\circ\mathbf{V})\boldsymbol{\alpha}=\boldsymbol{\sigma},$ (3)
where
$\mathbf{U}$ $=\begin{bmatrix}\mathbf{u}_{1}^{(1)\top}\!\mathbf{u}_{1}^{(1)}\!&\!\!\!\cdots\!\!\!&\mathbf{u}_{1}^{(1)\top}\!\mathbf{u}_{T}^{(r_{T})}\!\\ \vdots&\ddots&\vdots\\ \mathbf{u}_{T}^{(r_{T})\top}\!\mathbf{u}_{1}^{(1)}\!&\cdots&\mathbf{u}_{T}^{(r_{T})\top}\!\mathbf{u}_{T}^{(r_{T})}\end{bmatrix}$,
$\mathbf{V}$ $=\begin{bmatrix}\mathbf{v}_{1}^{(1)\top}\!\mathbf{v}_{1}^{(1)}&\!\!\!\cdots\!\!\!&\mathbf{v}_{T}^{(r_{T})\top}\!\mathbf{v}_{1}^{(1)}\\ \vdots&\ddots&\vdots\\ \mathbf{v}_{1}^{(1)\top}\!\mathbf{v}_{T}^{(r_{T})}&\cdots&\mathbf{v}_{T}^{(r_{T})\top}\!\mathbf{v}_{T}^{(r_{T})}\end{bmatrix}$,
$\boldsymbol{\alpha}$ $=[\alpha_{1}^{(1)},\cdots,\alpha_{T}^{(r_{T})}]\;\;\text{and}\;\;\boldsymbol{\sigma}=[\sigma_{1}^{(1)},\cdots,\sigma_{T}^{(r_{T})}].$